### v2

# Virtual states and exponential decay in small-scale dynamo

Abstract
We develop the Kazantsev theory of small-scale dynamo generation at small Prandtl numbers near the generation threshold and restore the concordance between the theory and numerical simulations: the theory predicted a power-law decay below the threshold, while simulations demonstrate exponential decay. We show that the exponential decay is temporary and owes its existence to the flattening of the velocity correlator at large scales. This effect corresponds to the existence of a long-living virtual level in the corresponding Schrodinger type equation.
We also find the critical Reynolds number and the increment of growth/decay above and under the threshold; we express them in terms of the quantitative characteristic properties of the velocity correlator, which makes it possible to compare the results with the data of different simulations.
Small-scale dynamo (SSD) in a turbulent flow is a widely discussed possible mechanism of magnetic field generation in various astrophysical objects Parker ; BSS-rev ; HazKit . The mechanism is based on stretching of magnetic lines trapped and carried by a small-scale turbulent flow, and implies high Reynolds numbers.
One differs two important special cases: the ratio of kinematic viscosity and magnetic diffusivity called magnetic Prandtl number
$Pm=\nu/\eta$
either much larger or much smaller than 1. The first case corresponds to, e.g., interstellar and intergalactic matter, while the second takes place in the Sun ($Pm\sim 10^{-6}\div 10^{-4}$; Stix ; Rempel ) and solar-type stars, and inside planets. The question whether the SSD plays an important role in solar magnetic field generation remains open, in particular because of significant difficulties in modeling low Prandtl numbers in experiments and simulations Sch07 ; Hotta . However, during last decade, direct numerical simulations were performed for magnetic Prandtl numbers down to $10^{-1}$ and even lower Sch07 ; I ; B ; W .
The theoretical investigations of small-scale dynamo are based on the Kazantsev approach Kaz . The Kazantsev equation relates the second order one-time magnetic field correlator and the longitudinal velocity correlator of some special type. For low Prandtl numbers, this equation was analyzed in, e.g., NRS ; VK ; Verg ; Rog97 ; Vinc ; Bold ; finny ; Malyshkin ; Kleeorin12 ; Schob ; Yu ; mnras ; the critical magnetic Reynolds number
Rm=PmReRm=Pm\,Re
corresponding to the threshold of generation was found NRS , and for $Rm>Rm_{c}$, the magnetic field growth rate near the threshold was shown to be Kleeorin12
$\gamma\sim\ln(Rm/Rm_{c})$ (1)
However, for $Rm$ below the generation threshold, from the Kazantsev theory it follows that the magnetic field perturbations spectrum is continuous; thus, the decay is not exponential, and $\left.\gamma\right|_{Rm<Rm_{c}}=0$. On the contrary, numerical simulations Sch07 ; I ; B ; W show that the linear dependence (1) still holds for $Rm<Rm_{c}$: in the graph $\gamma(\ln Rm)$ they present, the straight line continues under the abscissa axis.
In this paper, we eliminate this contradiction by means of an accurate analysis of Kazantsev equation in the region of scales comparable to the integral scale. It appears that, far as it is from the main region of magnetic field generation, it is still important. The Kazantsev equation can be written in the form of the Schrodinger equation with variable mass; the flattening of the kinetic energy spectrum at large scales results in some peculiar feature of this Schrodinger potential , namely, the presence of a positive-energy peak at these scales. The effect of the peak on the critical value of the magnetic Reynolds number is rather small, but it makes possible the existence of resonance positive-energy solutions: virtual levels. We argue that the observed exponential decay of the magnetic field correlator at $Rm$ right below the generation threshold is temporary, and corresponds to the virtual level. We show that $d\gamma/dRm$ is the same below and above the threshold, and estimate its value and the lifetime of the virtual level, which corresponds to the duration of the exponential damping.
I Kazantsev equation
The dynamics of the magnetic field advected by a turbulent flow is described by the equation
$\frac{\partial\mathbf{B}(\mathbf{r},t)}{\partial t}=\mathrm{rot}\bigl[\mathbf{v}(\mathbf{r},t)\times\mathbf{B}(\mathbf{r},t)\bigr]+\eta\Delta\mathbf{B}(\mathbf{r},t).$ (2)
Since the feedback of the magnetic field on the velocity dynamics is quadratic, and the seed magnetic field is assumed to be small, one can neglect the influence of magnetic field on the dynamics of the flow. At this (kinematic) stage of evolution the magnetic field is a passive vector, the velocity field ${\bf v}({\bf r},t)$ is considered stochastic with given stationary statistics; in Eq. (2) it acts as a multiplicative noise.
In this paper, following Kaz and many other papers, we consider Gaussian and delta time-correlated velocity statistics. Then, it is completely determined by the Lagrangian relative diffusivity of particles second order correlator:
$D_{ij}(\bm{\rho})=\int\langle v_{i}(\mathbf{r},t)v_{j}(\mathbf{r}+\bm{\rho},t+\tau)\rangle d\tau.$ (3)
We are interested in the pair correlation function $\left\langle B_{i}(\mathbf{r},t)B_{j}(\mathbf{r}^{\prime},t)\right\rangle$. To get the corresponding equation, one has to express its time derivative by means of (2), and then average over different realizations of the velocity field. The cross correlations of magnetic induction and velocity can be split by means of the Furutsu-Novikov theorem,
$\langle v_{p}(\mathbf{r},t)g[\mathbf{v}]\rangle=\frac{1}{2}\int D_{ij}(\mathbf{r}-\mathbf{r}^{\prime})\left\langle\frac{\delta g[\mathbf{v}]}{\delta v_{j}(\mathbf{r}^{\prime},t)}\right\rangle\mathrm{d}\mathbf{r}^{\prime},$ (4)
The conditions of isotropy and non-divergence of the velocity field reduce the tensor equation to only one component:
$G(\rho,t)={n}_{i}{n}_{j}\langle B_{i}(\mathbf{r}+\bm{\rho},t)B_{j}(\mathbf{r},t)\rangle$ (5)
Denote
$S(\rho)=\eta+\frac{1}{2}b(\rho)\ ,\quad b(\rho)={n}_{i}{n}_{j}\left(D_{ij}(0)-D_{ij}(\rho)\right)\ ,\quad\mathbf{n}=\bm{\rho}/\rho$
The function $b(\rho)$ is related to the second order velocity structure function by means of the expression:
$2b(\rho)=\int d\tau\left\langle\delta_{\parallel}\mathbf{v}(\bm{\rho},0)\delta_{\parallel}\mathbf{v}(\bm{\rho},\tau)\right\rangle\ \ ,\qquad\delta_{\parallel}\mathbf{v}(\rho,\tau)=(\mathbf{v}(\rho,\tau)-\mathbf{v}(0,\tau))\cdot{\bf n}\ \ ,$
where the index $\parallel$} means the longitudinal component.
The application of the Furutsu-Novikov formula produces the following equation :
$\frac{\partial}{\partial t}G(\rho,t)=2S(\rho)\left(G^{\prime\prime}_{\rho\rho}+\frac{4G^{\prime}_{\rho}}{\rho}\right)+2S^{\prime}G^{\prime}_{\rho}+2\left(S^{\prime\prime}+4\frac{S^{\prime}}{\rho}\right)G$ (6)
Following Kazantsev, we use the supposition
$\psi(\rho,t)=\rho^{2}\sqrt{S}\,G(\rho,t)\ ,$ (7)
and by means of the Laplace transform
$\psi=\int\limits_{0}^{\infty}e^{\gamma t}\psi_{\gamma}(\rho)d\gamma$ (8)
we arrive at the Schrodinger type equation
$\begin{array}[]{l}\displaystyle\psi^{\prime\prime}_{\rho\rho}=\frac{\gamma}{2S(\rho)}\psi+U(\rho)\psi\ ,\\ U=-\frac{1}{\rho^{2}}\left(\frac{3\sigma(\sigma+4)+1}{4}+\frac{\rho\sigma^{\prime}}{2}\right)\end{array}$ (9)
where
$\sigma(\rho)=\frac{\mathrm{d}\ln S}{\mathrm{d}\ln\rho}-1$
Here and below we omit the index $\gamma$ in $\psi_{\gamma}(x)$. In what follows we specify the explicit shape of the potential, and consider the existence of solutions to this equation. Formally, this equation is Schrodinger equation with variable mass; to proceed to the Sturm-Liouville problem, one has to make one more change of variables and to remove the dependence on $\rho$ in the term containing $\gamma$. For $Pr_{m}\gg 1$ this is done in (apj, , Appendix C). However, qualitative behavior of the solution in the case of variable mass remains the same, so here we analyze it without additional transformations.
In accordance with (7), we consider boundary conditions
$\psi(0)=0,\ \ \psi(\infty)<\infty$
We note that, since energy in Schrodinger equation corresponds to $\gamma$, the decreasing modes ($\gamma<0$}) correspond to the continuous spectrum, while for growing modes ($\gamma>0$) one obtains discrete spectrum; these discrete values of $\gamma$ produce the dynamo effect.
II Definition of parameters
In a turbulent flow, the structure function $b(\rho)$ demonstrates roughly piecewise power behavior; one can describe it by extending the Vainshtein-Kichatinov model VK :
$b\propto\left\{\begin{array}[]{ll}\rho^{2}\ ,&\rho\ll r_{\nu}\\ \rho^{1+s}\ ,&r_{\nu}\ll\rho\ll{\Lambda}\\ \rho^{0}\ ,&{\Lambda}\ll\rho\end{array}\right$.
Here the first line corresponds to the viscous range restricted by the Kolmogorov scale $r_{\nu}$, the second line describes the inertial range, and the third line demonstrates schematically the transition to the large-eddy scale where the power law breaks. Following Kazantsev, we leave the power $s$ undefined for now; the Kolmogorov theory predicts $s=1/3$ VK ; Rog97 , from intermittency it follows that $s>1/3$ Vain ; Frisch ; from the theoretical bridge relations LvovX combined with recent DNS results Iyer it follows that $s=0.39$ for asymptotically large Reynolds numbers.
The graph of this function in logarithmic coordinates has three linear sections; in real turbulence, of course, they are separated by two transition zones. The first transition zone between viscous and inertial ranges is called bottleneck; it does not affect generation seriously since for very small magnetic Prandtl numbers the magnetic diffusivity scale lies deep inside the inertial range. So, we are interested in the scales $\rho\gg r_{\nu}$, and thus can neglect the effects of viscosity and the bottleneck (i.e., the transition from viscous to inertial range). Instead, we concentrate on the inertial range and the transition to the integral scale. The second transition zone is essential for magnetic field generation paper140 , but to understand the main idea of the effect we do not have to consider it accurately.
To make the possibility of comparison of experiments and simulations with theory predictions more transparent, we now propose formal and model-independent definitions to the main properties of $b(\rho)$ that affect the generation.
Let $b_{\infty}=b(\rho\to\infty)$ be the largest-scale limit of $b$; then we define
$X=\left(b_{\infty}/2\eta\right)^{1/(1+s)}$
The scale $r_{d}$ is the one where molecular diffusion balances turbulent diffusion. To define it accurately, we choose
$r_{d}\,:\ \ \frac{1}{2}b(r_{d})=\eta$
The value of $X$ can be interpreted as the width of the ’turbulent generation range’. In the frame of the model, we get $\Lambda=r_{d}X$. However, in real turbulence the derivative of $b$ changes smoothly, there is no sharp transition, and the value of $\Lambda$ is not well defined. To the contrary, the values of $r_{d}$ and $X$ are universally defined and can be measured for any Reynolds number and for any geometry of a flow.
The magnetic Reynolds number can be approximately calculated by the following consideration:
$Rm=\frac{LV}{\eta}\sim\frac{v_{rms}^{2}T}{\eta}\$
where $V$ is the characteristic velocity scale and $L,T$ are the correlation length and time of the largest-scale vortices; on the other hand, from the definition of the correlation time it follows
$b_{\infty}=\left.\int d\tau\left\langle\delta v_{\parallel}(\rho,0)\delta v_{\parallel}(\rho,\tau)\right\rangle\right|_{\rho\to\infty}=\langle(\delta_{\parallel}\,v(\infty))^{2}\rangle T\simeq\frac{2}{3}v_{rms}^{2}T$
Thus, we get
$Rm\simeq\frac{3b_{\infty}}{2\eta}=3\cdot X^{1+s}$
However, this is only a rough estimate to the order of magnitude, since different simulations use different definitions of $L$. For more accurate comparison of $X$ and $Rm$ , one has to investigate the properties of $b(r)$ used in every particular simulation. An attempt of such an accurate analysis will be made in our paper paper140 .
III The shape of the potential
We now pass on to the dimensionless variable
$x=\rho/r_{d}$
The function $S$ then takes the form:
$S=\eta\tilde{S}(x)=\eta\left(1+x^{1+s}\right)\theta(X-x)+\eta\left(1+X^{1+s}\right)\theta(x-X)$
The corresponding expression for the logarithmic derivative is
$\sigma+1=(1+s)\frac{x^{1+s}}{1+x^{1+s}}\theta(X-x)$ (10)
From (9) we get
$\psi^{\prime\prime}_{xx}=\frac{\tilde{\gamma}}{\tilde{S}(x)}\psi+\tilde{U}(x)\psi\ ,\quad\tilde{U}=-\frac{1}{x^{2}}\left(\frac{3\sigma(\sigma+4)+1}{4}+\frac{x\sigma^{\prime}_{x}}{2}\right)$ (11)
where
$\tilde{\gamma}=\frac{r_{d}^{2}}{2\eta}\gamma$ (12)
We pay special attention to the fact that not only $\sigma$ but also its derivative is present in the potential. The step change in $\sigma$ at $x=X$ thus results in a $\delta$-function in the potential (see Fig.1). This implies a nontrivial matching condition for the solutions at the point $X$: the spatial derivative of the magnetic field correlator has a break.
Figure 1: The shape of the potential for $s=1/3$. The $\delta$-function is an artifact of the model, it does not affect the solutions essentially. The presence of a ’thick’ maximum is of decisive importance.
We consider the range of parameters that corresponds to the boundary of generation; this happens at large magnetic Reynolds numbers, which corresponds to large values of $X$. For $X\gg 1$, (10) simplifies to
$\sigma+1\simeq(1+s)\theta(X-x)$
and the matching condition takes the form
$\psi_{<}(X)=\psi_{>}(X)\ ,\quad X\frac{\psi^{\prime}_{>}(X)}{\psi(X)}-X\frac{\psi^{\prime}_{<}(X)}{\psi(X)}=\frac{1+s}{2}$ (13)
Thus, one has to solve Eq. (11) with account of the matching conditions (13).
IV Critical Reynolds number
To find the boundary of the generation, we set $\gamma=0$ in (11). To the left of the boundary, $1\ll x<X$ , the asymptotic expression for the potential is
$\tilde{U}=-\frac{1}{x^{2}}\left(\kappa^{2}+\frac{1}{4}+O(x^{-(1+s)})\right)$ (14)
where
$\kappa=\sqrt{\frac{3s(s+4)}{4}}$
The corresponding asymptote of the solution is
$\psi_{<}\propto\sqrt{x}\cos\left(\kappa\ln x+\varphi\right)\ ,\ \ 1\ll x<X$ (15)
To the right of the boundary, the physically reasonable normalizable solution is $\psi_{>}\propto 1/x$. The matching condition (13) becomes
$\kappa\tan\left(\kappa\ln X+\varphi\right)=2+\frac{s}{2}$ (16)
The phase $\phi$ is determined by the behavior of the solution at small $x$. To find it, one has to consider the exact solution of (11) for $x<X$ that converges at small $x$ :
$\psi_{<}\propto x^{2}$ $\sqrt{1+x^{1+s}}\,\,{}_{2}F_{1}\left(\alpha-\mathrm{i}\,\beta,\alpha+\mathrm{i}\,\beta,2\alpha,-x^{1+s}\right)$ (17)
$\,\,\,\,\,\,\,\,\,\alpha(s)=\frac{4+s}{2(1+s)}\,\,\,\,\,\,\beta(s)=\frac{\kappa}{(1+s)}$
Comparing this with (15) and making use of
$_{2}F_{1}\left(a,b,c,-z\right)=\frac{\Gamma(c)\Gamma(b-a)}{\Gamma(b)\Gamma(c-a)}\,z^{-a}+\frac{\Gamma(c)\Gamma(a-b)}{\Gamma(a)\Gamma(c-b)}\,z^{-b}\ ,\ \ \ z\gg 1\ ,$
one can find
$\varphi=\mathrm{Arg}\left\{\frac{\Gamma\bigl(2\mathrm{i}\beta\bigr)}{\Gamma^{2}\bigl(\alpha+\mathrm{i}\beta\bigr)}\right\}.$ (18)
The smallest positive solution of Eq. (16 ) , (18) for $s=1/3$ is
$X_{c}=20.5$
This is the value of $X$ that corresponds to the generation threshold; other solutions correspond to appearance of next levels in the potential well. The corresponding critical value of $Rm$ is
$Rm_{c}\simeq 3X_{c}^{4/3}\sim 100$
This estimate depends significantly on the exact definition of the Reynolds number. For detailed comparison of this theoretical result with the results of numerical simulations, see paper140 .
Without the last term in the potential, the matching conditions would take the form $\psi_{<}=\psi_{>}$, $\psi^{\prime}_{<}=\psi^{\prime}_{>}$, and one would get $\kappa\tan\left(\kappa\ln X+\varphi\right)=3/2$ instead of (16). This corresponds to the potential obtained in Kleeorin12 ; Schob and results in $X_{c}({\rm without\,\delta})=17.6$, so the value of $Rm_{c}$ calculated without the account of $\delta$-function in the last term is underestimated to about 20% .
V Below and over the critical Reynolds number
Now we consider small but non-zero $\gamma$. In the region $x\gg 1$, we have asympthotic expressions for Eq. (11) in the two zones:
$\psi^{\prime\prime}_{xx}$ $=$ $\frac{\tilde{\gamma}}{x^{1+s}}\psi-\frac{\kappa^{2}+1/4}{x^{2}}\psi\ \ ,\ x<X$ (19)
$\psi^{\prime\prime}_{xx}$ $=$ $\frac{\tilde{\gamma}}{X^{1+s}}\psi+\frac{2}{x^{2}}\psi\ \ ,\ \ \ \ x>X$ (20)
The exact solutions to (20) can be expressed in terms of Bessel functions of imaginary argument. The approximation for small $\tilde{\gamma}$ can be easily found from (19):
$\psi_{<}(x)=A\sqrt{x}\left[\cos a+\frac{\tilde{\gamma}x^{1-s}}{(1-s)^{2}}\cos p\cos\left(a-p\right)\right]\ ,$ (21)
$a(x)=\kappa\ln x+\varphi\ ,\ \ p=\arctan(2\kappa/(1-s))$
Here we assume that the phase $\varphi$ coincides with that found in (18) for $\gamma=0$. Actually, for small $x$ $\tilde{U}(x)\gg\tilde{\gamma}/\tilde{S}$ , and the influence of $\gamma$ on $\varphi$ is small.
V.1 Generation: $\gamma>0$
To the right from $X$, the general solution to (20) is
$\psi_{>}(x)\propto\left(\frac{1}{x}+\sqrt{\tilde{\gamma}X^{-1-s}}\right)\exp(-\sqrt{\tilde{\gamma}X^{-1-s}}\,x)$
We are interested only in the nearest vicinity of the boundary $X$, and in the smallest $\tilde{\gamma}$; For $\tilde{\gamma}\ll X^{s-1}$ and $x\gtrsim X$ this can be simplified to the first-order approximation in $\tilde{\gamma}$ :
$\psi_{>}(x)=B\left(\frac{1}{x}-\frac{\tilde{\gamma}}{2X^{1+s}}x\right)$ (22)
So, it is convenient to present the matching conditions (13) in the form
$\begin{array}[]{l}\psi_{<}(X)=\psi_{>}(X)=\frac{B}{X}(1-\frac{1}{2}\tilde{\gamma}X^{1-s})\ ,\\ \psi+X\psi_{<}^{\prime}+\frac{1+s}{2}\psi=\psi+X\psi_{>}^{\prime}=-\frac{B}{X}\tilde{\gamma}X^{1-s}=-\tilde{\gamma}X^{1-s}\psi+O((\tilde{\gamma}X^{1-s})^{2})\end{array}$ (23)
Substituting (21) in the second equality, we get
$\kappa\sin a-(2+\frac{s}{2})\cos a=\tilde{\gamma}X^{1-s}\left[\frac{\cos p}{(1-s)^{2}}\left((3-\frac{s}{2})\cos(a-p)-\kappa\sin(a-p)\right)+\cos a\right]$
Here $a$ is taken at the point $X$, i.e., $a=a(X)$. This equation determines the deviation of $X$ and the corresponding $a$ from the critical values $X_{c}$ and $a_{c}$ for given $\tilde{\gamma}$. In the limit $\tilde{\gamma}=0$ , the right-hand side is zero, and we arrive at (16). Thus, for small $\tilde{\gamma}$ the value of $a$ is close to $a_{c}$, and $\delta a=a-a_{c}\propto\tilde{\gamma}X^{1-s}$. Neglecting the second order in $\tilde{\gamma}$ and taking into account the substitutions for $p$ and $a_{c}$, we obtain the coefficient:
$\frac{da}{d\tilde{\gamma}X_{c}^{1-s}}=\frac{4\kappa}{4\kappa^{2}+(4+s)^{2}}\left(1+\frac{2\kappa^{2}+(4+s)(3-\frac{s}{2})+(1-s)^{2}}{(1-s)((1-s)^{2}+4\kappa^{2})}\right)$
Finally, the substitution of the explicit form of $\kappa$ results in
$\tilde{\gamma}=\ln\frac{X}{X_{c}}X_{c}^{s-1}\cdot(4+s)\frac{(1-s)(4s^{2}+10s+1)}{2(7-2s^{2})}+O(\ln^{2}(X/X_{c}))$ (24)
Restoring the dimensional parameters by means of (12) and taking into account that $Rm\propto X^{1+s}$ , we finally arrive at
${\gamma}=\ln\frac{Rm}{Rm_{c}}\frac{cX_{c}^{s-1}}{1+s}\frac{2\eta}{r_{d}^{2}}$ (25)
where
$c=(4+s)\frac{(1-s)(4s^{2}+10s+1)}{2(7-2s^{2})}$
By chance, for $s=1/3$, this coefficient is to good accuracy equal to 1. So, we obtain the coefficient in the logarithmic dependence of $\gamma$ on $Rm$. By means of the method proposed in (blobs, , Appendix A), we confirm this result from Eq. (11) numerically. To obtain the coefficient from DNS one needs the value of $r_{d}$. For accurate analysis of data of different papers and for their comparison with this result, we refer the reader to paper140 . Here we restrict ourselves by very rough estimate to the order of magnitude, to make sure that the theory does not contradict experimental/simulation data.
In B ; W the corresponding relation is presented in the form
$\gamma_{DNS}=0.022v_{rms}k_{f}\ln(Rm/Rm_{c})$
To compare this with our result, we note that $\Lambda=Xr_{d}$ is of the same order as the integral scale $L=2\pi/k_{f}$ . Taking into account $Rm=Lv_{rms}/\eta\simeq 3X^{1+s}$ and keeping in mind that everything happens near the generation boundary, at $Rm=Rm_{c}$, we get
$\frac{d\gamma}{d\ln(Rm/Rm_{c})}=\frac{cX_{c}^{s-1}}{1+s}\frac{2\eta}{r_{d}^{2}}\sim\frac{cX_{c}^{s-1}}{1+s}\frac{2X^{2}v_{rms}}{LRm}\sim\frac{c}{3\pi(1+s)}k_{f}v_{rms}$
This estimate gives the coefficient $0.08k_{f}v_{rms}$ , which does not differ radically from the results of the simulations.
V.2 Damping: $\gamma<0$
Conside some fixed $X<X_{c}$, $X_{c}-X\ll X_{c}$. There are no solutions with positive $\gamma$, so the magnetic field correlations decay. The solution to Eq. (11) to the right from $X$ ($x>X$) is now described by a two-parametric oscillating function, so one can match the left-hand and right-hand sides for any $\gamma$. In terms of quantum mechanics, this continuous spectrum corresponds to positive energy of free particles. Since solutions with all $\gamma<0$ are present, the decay is slower than any exponential. (Also in two-dimensional turbulence, the decay is always power-law Kol ).
Thus, the Kazantsev theory predicts power-law decay for magnetic Reynolds numbers under the generation threshold. However, it turns out that just below the threshold, for each $X$ there is an exponent $\gamma(X)$ such that its amplitude in the inner region is much higher than the amplitudes corresponding to other exponents. This effect is analogous to a virtual level in quantum mechanics, and results in existence of temporal exponential damping with the exponent $\gamma(X)$. After some time, the exponential behavior changes back to a power-law.
For brevity, we now introduce one more parameter. Denote
$q=-\tilde{\gamma}/X^{1+s}>0$
Then the solution to Eq. (20) in the region $x>X$ is
$\psi_{>}=B\left(\frac{\sin(x\sqrt{q}+\delta)}{x\sqrt{q}}-\cos(x\sqrt{q}+\delta)\right)$ (26)
where $B,\delta$ are constants. The condition (13) now establishes the matching between (26) and (21).
We note that from (26) it follows
$\begin{array}[]{l}f(x)=\psi_{>}+x\psi_{>}^{\prime}=Bx\sqrt{q}\sin(x\sqrt{q}+\delta)\ ,\\ f(x)-x^{2}q\,\psi_{>}=Bx^{2}q\cos(x\sqrt{q}+\delta)\end{array}$ (27)
To exclude $\delta$, we consider the combination
$x^{2}qf^{2}+\left(f-x^{2}q\,\psi_{>}\right)^{2}=B^{2}x^{4}q^{2}$
and apply it at the matching point:
$B^{2}X^{4}q^{2}=f^{2}(X)X^{2}q+\left(f(X)-X^{2}q\,\psi_{>}(X)\right)^{2}$ (28)
What is known about the ’outer’ amplitude $B$? It is, generally, a function of $q$ determined by the expansion of the initial function $\psi(t,x)$ in a series of the eigenfunctions (26) with different $q$. Assume that (unlike the case considered in Yu ) the initial function $\psi(t=0,x)$ is a function with a compact support $\ll X$. Then, according to (8), $\int\psi(x,q)dq=0$ for any $x$ larger than some $x_{1}$; the same is true for $f=\psi_{>}+x\psi_{>}^{\prime}$. From (27) we then get
$\int Bx\sqrt{q}\sin(x\sqrt{q}+\delta)dq=\int 2Bx{q}\sin(x\sqrt{q}+\delta)d\sqrt{q}=0\ ,\ \ x>x_{1}$
and from the inverse Fourier transform it follows that $B\propto 1/q$ for small $q$. So, the left-hand side of (28) is nearly constant for small $q$.
By means of (13), one can express the right-hand side of (28) via $\psi_{<}$, $\psi^{\prime}_{<}$, and then substitute (21). Then (28) represents the proportionality between the ’inner’ and ’outer’ amplitudes $A$ and $B$. To select the value of $q$ with the largest $A$, one has to find the minimum of the proportionality coefficient on the right-hand side.
We extract $A$ explicitly from the right-hand side of (28) and expand the rest into a series in $q$; Eq. (21) confirms the correctness of this procedure:
$f(X)=A\tilde{f}=A(f_{0}+qf_{1}+...)\ ,\ \ \ \psi(X)=A\tilde{\psi}=A(\psi_{0}+q\psi_{1}+...)\ ,\ \ B^{2}X^{4}q^{2}=A^{2}F(q)$
We also note that, because of (16), $f_{0}(X_{c})=0$ and $f_{0}(X)\propto X-X_{c}$ is small. Thus, $f_{0}\ll f_{1}\sim\psi_{0}\sim f_{2}\sim$... Then, taking the derivative of $F$ with respect to $q$ and setting it zero, we find to the first order in $q$:
$f(X)=X^{2}q\,\psi(X)+O(q^{2})$ (29)
This is the condition that singles out the $q(X)$ corresponding to the maximal amplitude $A$. Since the first summand in (28) is parametrically small as compared to the second, it is natural that the requirement is fulfilled if the second summand is equal to zero.
This expression estabishes the correspondence between the value of $X$ and the the largest -amplitiude mode $\gamma$ . This is the quasi-level: for this $X$, the contribution to $\psi$ that decays with the selected $\gamma$ is (temporarily) stronger than all others.
[…]
Width of the virtual level. The second derivative of $B^{2}X^{4}q^{2}/A^{2}$ is
$d^{2}F/dq^{2}=2(f_{1}-X^{2}\psi_{0})^{2}+O(q)$
The value $A^{2}$ is two times smaller at the point where the factor $F$ is two times larger than at the minimum, i.e.,
$\frac{1}{2}F^{\prime\prime}(q_{min})\Delta q^{2}=F(q_{min})=qX^{2}f^{2}$
So, the relative width of the quasi-level is approximately
$\frac{\Delta\gamma}{|\gamma|}=\frac{\Delta q}{q}=\frac{\sqrt{q}Xf}{q|X^{2}\psi_{0}-f_{1}|}\simeq\frac{q^{1/2}X^{3}\psi}{|X^{2}\psi-f_{1}|}\sim q^{1/2}X\sim|\tilde{\gamma}|^{1/2}X^{(1-s)/2}$ (30)
This confirms that the virtual level is rather narrow for small $\gamma$, and its lifetime exceeds essentially the characteristic timescale of the decay.
Comparison of (30) with (24) shows that the virtual level exist while $\ln(X_{c}/X)\lesssim 1$.
Thus, for $Rm$ less than, say, one half of $Rm_{c}$ the intermediate exponential asymptote is rather short in time, and for smaller $Rm$ the virtual level vanishes.
This estimate of the minimal $Rm$ that implies a virtual level is in concordance with the one that can be obtained from the height of the potential right after the peak (the delta function itself does not contribute to the appearance of virtual levels, only the ’thick’ part of the peak matters). Assuming that the boundary ’energy level’ equals to this height, we get from (11)
$\tilde{\gamma}\tilde{S}\lesssim U(X+0)=2/X^{2}\ \ \rightarrow-\tilde{\gamma}\lesssim X^{1+s}/X^{2}\sim X^{s-1}$ (31)
which, to the coefficient $\sim 1$, coincides with the $\tilde{\gamma}$ obtained from (24) for $\ln(X/X_{c})\lesssim 1$.
The lifetime of the level for small $|\gamma|$ can be estimated by
$t\sim\frac{1}{\Delta\gamma}\sim\frac{\sqrt{\eta}}{r_{d}\,|\gamma|^{3/2}\,X^{(1-s)/2}}$ (32)
Denote
$T=\frac{r_{d}^{2}}{\eta}X^{1-s}$
Then (30) is written as
$\frac{\Delta\gamma}{|\gamma|}\sim\sqrt{|\gamma|\,T}$ (33)
and (32) becomes
$t\sim\frac{1}{T^{1/2}\,|\gamma|^{3/2}}$ (34)
The value of $T$ is equal to $T=\Lambda^{2}/b_{\infty}\sim L/v_{rms}$; generally, it corresponds to the characteristic largest-eddies turnover time.
VI Conclusion
In the paper we consider the evolution of the magnetic field passively advected in a turbulent flow. We study the flows with high Reynolds and small magnetic Prandtl numbers near the generation threshold, and use the Kazantsev theory with the Vainshtein-Kichatinov model for the turbulent velocity statistics.
We examine the kink and the peak in the Kazantsev potential at the scales corresponding to the transition region from the inertial range to the largest-eddies scales. These features of the potential are a consequence of the flattening of the velocity structure function at these scales. We show that these details of the potential produce a virtual energy level, which corresponds to the virtual state with temporary exponential decay for magnetic Reynolds numbers below the generation threshold.
Based on the Kazantsev theory, we show that the linear dependence $\gamma\propto\ln Rm/Rm_{c}$ continues also for negative argument, in full agreement with the results of the DNS Sch07 ; I ; B ; W . We predict that in longer observation, the exponential decay will stop after some time and be replaced by a slower decrease. We give an estimate of this time (34) and show that it is asymptotically long for magnetic Reynolds numbers close to the generation threshold, and by the end of the exponential stage the magnetic field correlator would decrease by $\exp{|\gamma|/\Delta\gamma}\sim\exp{|\gamma|^{-1/2}}$ times (33). For Reynolds magnetic numbers far from the critical value, approximately $Rm\lesssim Rm_{c}/2$, there is no stage of exponential decay.
Although the results were obtained in the frame of the extended Vainstein-Kichatinov model, the existence of the virtual level and, accordingly, the temporal exponential decay of the magnetic field correlator just below the generation threshold is a consequence of the flattening of the velocity correlation function at large scales, and hence, is a general feature of all reasonable models. It also seems that finite correlation time of the velocity field, its time irreversibility or helical asymmetry would not affect the result.
The work of AVK was supported by the RSF grant 24-72-00068.